OpenStudy (anonymous):

if \[\sin \frac{ x }{ 2 }=\frac{ 2 }{ 3 }, \left[ 0,\frac{ \pi }{ 2 } \right]\] Evaluate. sin(x/4)

OpenStudy (freckles):

hey

OpenStudy (freckles):

just so you know 1/2*x/2=x/4

OpenStudy (freckles):

makes me think we need half angle identity \[\sin^2(\frac{1}{2} \cdot u)=\frac{1}{2}(1-\cos(u)) \\ \text{ where our } u=\frac{x}{2}\]

OpenStudy (freckles):

you are given sin(x/2)=2/3 where x is between 0 and pi/2 draw a right triangle in the first quadrant and find cos(x/2) because you will need this value

OpenStudy (anonymous):

I don't quite understand why we use cosine for this question..

OpenStudy (freckles):

do you see the half angle identity above?

OpenStudy (freckles):